(2) Determine k such that kx^2 - 4x + 1/2=0 has real and different solutions. (3) (i) What rate of interest compounded annually is needed to make an investment of $6,500 accumulate to an amount of $7166.25 at the end of two years? (ii) How long will it take for the initial amount to double, compounded annually at the rate in (i) above?


1
Expert's answer
2021-11-11T14:38:55-0500

2)

The polynomial equation is,ย "kx\\^2-4x+\frac{1}{2}=0"

Where "a = k, b= -4, c=\frac{1}{2}"

Since the equation has distinct root;

therefore "b\\^2-4ac > 0"

"(-4)\\^2-4\times k \times\frac{1}{2} > 0\\
16 - 2k>0\\
16>2k\\
8>k" , having two real roots,.


3)

i)

Compounded annually formula is "A = P(1 + r)\\^t"

A = $7166.25

t = 2

P $6500

Therefore,

"7166.25 = 6500(1 + r)\\^2\\
\frac{7166.25}{6500} = (1 + r)\\^2\\
1.1025 = (1 + r)\\^2\\

\\\sqrt{1.2025} - 1 = r\\
r = 0.050013"

Convert rate of interest as percentage: "R = r \times 100\\"

Rates (R) = 5.002 per year.

ii) The amount double = "2 \times 6500 = \$13000\\"

A = $13000

Principal = $6,500

Rate of interest = 5.002%

t= t

"13000 = 6500(1 + 0.05002)\\ ^ t\\
2 = (1.05002)\\ ^ t"

Taking log on both sides:

"\log2 = t\log(1.05002)\\
t = \frac{\log 2}{\log 1.05002}\\
14.201 = t"

Time for amount to be doubled is 14 years and 2 months